OpenStudy (anonymous):

Will Give Medals. An equation is shown below: 8(2x - 14) + 13 = 4x - 27 Part A: Write the steps you will use to solve the equation, and explain each step. Part B: What value of x makes the equation true?

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

First step is to distribute the 8 to 2x-14

OpenStudy (anonymous):

Ok

OpenStudy (sleepyhead314):

Hello @TheWebAcademyStudent and Welcome to OpenStudy! :) 8(2x - 14) + 13 = 4x - 27 first we should distribute the 8 and get 16x - 112 + 13 = 4x - 27 then add like terms 16x - 99 = 4x - 27

OpenStudy (anonymous):

Um Ok

OpenStudy (anonymous):

Thanks!

OpenStudy (sleepyhead314):

then to get all the x's to one side, you should subtract 4x from both sides 16x - 99 = 4x -27 - 4x -4x ------------------- 12x - 99 = -27 then add 99 to both sides +99 +99 ----------------- 12x = -27 + 99

OpenStudy (anonymous):

Ohh ok

OpenStudy (anonymous):

1) Distribute 8 16x-112+13=4x-27 2) Isolate x to one side by moving the lower variable (4x) Subtract 4x from each side: 12x-112+13= -27 3)Isolate x again Subtract 13 from each side: 12x-112=-40 Add 112 to each side: 12x=72 Divide 12 from each side: x=6